OpenStudy (mendicant_bias):

(Sturm-Liouville) I'm having trouble in dealing with SLDE's in determining the Eigenfunction. I can generally determine the eigenvalues without too much struggle, but I am failing to understand how to systematically arrive at eigenfunctions for a given eigenvalue. More info/example below.

OpenStudy (mendicant_bias):

In this document in Problem 1 (attempting to work through the problem before looking at solution) I was able to get all the way to the Eigenvalue, but I don't understand why the eigenfunction is a sine function. http://www2.math.uu.se/~gaidash/1MA208/Files/Solutions_2.pdf @Kainui , could you help me out with this?

OpenStudy (mendicant_bias):

(Or @freckles )

OpenStudy (freckles):

do you understand how they got their u?

OpenStudy (freckles):

\[u=A \cos(w \ln(x))+B \sin(w \ln(x)) ? \\ \text{ we apply conditions } \\ \text{ first I will find } u' \text{ before doing so } \\ u'=\frac{-Aw}{x}\sin(w \ln(x))+\frac{Bw}{x} \cos(w \ln(x)) \\ \text{ anyways } u(1)=0 \text{ means we have } \\ 0=A \cos(0)+Bsin(0)=A+0 \\ \text{ which means } A=0 \\ \text{ so far we have } u=B \sin(w \ln(x)) \text{ as the solution }\]

OpenStudy (freckles):

we need to apply the other condition

OpenStudy (freckles):

but cos should definitely not be apart of the egienfunction there since A=0

OpenStudy (freckles):

now let's look at the u' think and we only need to look at u'=Bw/x * cos(w ln(x)) since A=0

OpenStudy (freckles):

\[u'(e)=\frac{Bw}{e}\cos(w)=0\]

OpenStudy (freckles):

this means either B=0 or w=0 or cos(w)=0 We don't want B=0 and we don't want w=0 B=0 means we have a trivial solution and we looking for a nontrivial solution w=0 means we will also have a trivial solution since it will mean lambda is 0 then we would have (xu')'=0 which implies xu'=c u'=c/x u=c ln|x|+d and using the condition u'(e)=0 we have 0=c/e which means c is 0 anyways so then you have u=d but d is also 0 since u(1)=0 So that means the only thing that we could allow to be 0 is the cos(w) thing since this would lead to a nontrivial solution

OpenStudy (freckles):

@Mendicant_Bias

OpenStudy (mendicant_bias):

Sorry, was in lecture. I'm looking at this now.

OpenStudy (freckles):

ok i will try to clarify anything I said above if you need more clarification on anything

OpenStudy (mendicant_bias):

Heh, I think I linked the wrong document, but this is nonetheless helpful and may be nearly the exact same problem

OpenStudy (mendicant_bias):

Yeah, the one I intended to link was this one: http://www.ima.umn.edu/~miller/4567midtermsolns2S10.pdf

OpenStudy (mendicant_bias):

(Problem 3)

OpenStudy (mendicant_bias):

@freckles

OpenStudy (mendicant_bias):

So, no Cauchy-Euler, no natural logs involved in the complex roots solution, and slightly different Boundary Conditions. I evaluate it like this: (one second)

OpenStudy (mendicant_bias):

\[y''+\lambda y=0, \ \ \ 0<x<\pi, \ \ \ y(0)=0, y'(\pi)=0.\]\[y_h=c_1\cos(\alpha x)+c_2\sin(\alpha x) \]

OpenStudy (mendicant_bias):

\[y(0)=c_1\cos(\alpha x) +c_2 \sin(\alpha x)=0, \]\[c_1\cos(0)=0, \ \ \ \mathbf{c_1=0}\]

OpenStudy (mendicant_bias):

\[y'(\pi)=[-c_1 \sin(\alpha x)+c_2\cos(\alpha x)]\alpha=0\]

OpenStudy (freckles):

\[y(0)=c_1(\cos( \alpha \cdot 0))+c_2 (\sin( \alpha \cdot 0))=c_1+0 \\ y(0)=c_1 \\ \text{ but since } y(0)=0 \text{ then } c_1=0 \\ \text{ so far we can only have this as a possible nontrivial solution } y=c_2 \sin(ax )\]

OpenStudy (mendicant_bias):

\[c_2 \alpha \cos(\alpha \pi)=0, \ \ \ \alpha \pi=\frac{(2n-1)\pi}{2}, \ \ \ \alpha =\frac{(2n-1)}{2}\]

OpenStudy (mendicant_bias):

(Solving for eigenvalue where lambda = alpha squared, etc etc, just confused about how to find the eigenfunction for this case)

OpenStudy (freckles):

the eigenfunction is the solution that doesn't give us a trivial case

OpenStudy (freckles):

the trivial case being y(x)=0

OpenStudy (mendicant_bias):

(One minute, trying to phrase my questions correctly)

OpenStudy (freckles):

\[y=c_2 \sin(\alpha x) \\ y'=c_2 \cdot \alpha \cos(\alpha x) \\ y'(\pi)=c_2 \alpha \cos(\alpha \cdot \pi) \\ \text{ but } y'(\pi)=0 \text{ so this means we have } \\ c_2 \alpha \cos( \alpha \cdot \pi)=0 \\ \text{ case 1: } c_2=0: \text{ gives } y=0+0=0 \text{ this is a trivial solution } \\ \text{ we do not want to consider that possibility } \\ case 2: \alpha=0: \text{ gives } \lambda=0 \text{ since } \lambda=\alpha^2 \text{ which means you would have the diff equa } \\ y''(x)=0 \text{ with the conditions you gave } y'=c \\ y''=c+dx \\ y'=c \\ y'(\pi)=0 \text{ tells us } c=0 \\ y(0)=0 \text{ tells us } d =0 \\ \text{ so } \alpha=0 \text{ gives a trivial solution \because } y=0+0 =0 \text{ again } \] so the only other thing that can give us a possible nontrivial solution and also satisfy the equation: \[c_2 \alpha \cos( \alpha \cdot \pi)=0 \\ \text{ is allowing } \cos( \alpha \cdot \pi)=0 \\ \text{ is the third case } \\ \text{ if } \cos(\alpha \cdot \pi)=0 \text{ then } \alpha \cdot \pi=n \cdot \frac{\pi}{2} \text{ where } n \text{ is an add integer } \\ \alpha \cdot \pi=\frac{n \pi}{2} \text{ implies } \alpha=\frac{ n }{2} \text{ where } n \text{ is odd } \\ \text{ so recall that we had } y=c_2 \sin(\alpha x) \text{ and we just found from the second } \\ \text{ condition that } \alpha \text{ needs to be } \frac{ n}{2} \text{ where again } n \text{ is odd }\]

OpenStudy (freckles):

add=odd i'm done editing that

OpenStudy (mendicant_bias):

Their value for alpha is different than yours, and so is mine if I've read correctly, I'm a little confused as to why alpha isn't \[\alpha=\frac{(2n-1)}{2}\]

OpenStudy (freckles):

actually the alpha is not different

OpenStudy (freckles):

my is n/2 where is odd integer their's (2k-1)/2 where k is an integer 2k-1 is add for any integer k

OpenStudy (freckles):

omg odd*

OpenStudy (mendicant_bias):

Ah, just missed where you wrote that n is odd, got it

OpenStudy (freckles):

probably because i called it add integer :p

OpenStudy (freckles):

but I think I also referred to as odd integer too in the same paragraph

OpenStudy (mendicant_bias):

Alright, so procedurally, you first solve for your eigenvalues after finding the general solution and applying boundary conditions, and then, *after* you have your eigenvalues that work, you plug them back into the general solution along with boundary conditions in order to get your eigenfunction.

OpenStudy (freckles):

yeah i see which of the possible egienvalues gives me eigenfunctions by pluggin in to the orginial if I don't get y=0 then the possible eigenvalue is an actual eigenvalue and whatever y equals from that is my eigenfuction

OpenStudy (freckles):

and I didn't do case 3

OpenStudy (freckles):

their case 3

OpenStudy (freckles):

in case 3 they used their conditions for the general from and seen that A=B=0 for lambda in the form of -alpha^2 that is trivial solution having y=0 so there are no egienvalues and therefore no egienfunctions

OpenStudy (mendicant_bias):

Yep, got that part. Alright, awesome, thank you so much, I needed clarity on this.

OpenStudy (freckles):

Ok so you are good, for real? like there were no other possibilities for case 3 e^(u)+e^(v) is never 0 for any v or u because e^u>0 for all u and e^v>0 for all v so that means that we can only have A=0 and since B=-A then B=0 which gives us y=0

OpenStudy (mendicant_bias):

Yeah, I'm aware, the others made sense to me due to them being trivial, I literally just did not know the proper procedure to find an eigenfunction (but understood how to find eigenvalues).